OpenStudy (needhelpp101):

Solve the problem. During the 1998-1999 Junior Hockey League season, the Sharks played 54 games. Together, their wins and loses totaled 46. They tied 5 fewer games than they lost. How many games did they win that season? A. 8 games B. 33 games C. 13 games D. 29 games

OpenStudy (anonymous):

Lets set up some variables w +l = 46 t = l-5 w+l+t = 54

OpenStudy (anonymous):

by interchanging the first equation into the last one - you should be able to determine value of T

OpenStudy (anonymous):

from that - determine value of L with 2nd equation and last, determine W value with 1st equation. Please tell me what you got for T L and W

OpenStudy (matt101):

Say W is the number of wins, L is losses, and T are tie games. You know that W + L + T = 54 from the first sentence You know that W + L = 46 from the second sentence (meaning T = 8). You know that T = L - 5 from the third sentence (so 8 = L - 5 --> L = 13) Now that we have both T and L, you can easily find W from the above :)